In the following experiment, a coffee-cup calorimeter containing 100 mL of H2O is used. The initial temperature of the calorimeter is 23.0 ∘C. If 5.30 g of CaCl2 is added to the calorimeter, what will be the final temperature of the solution in the calorimeter? The heat of solution ΔHsoln of CaCl2 is −82.8 kJ/mol
Answer #1

Supposing the calorimeter itself absorbs no heat:

Mw CaCl2= 110.984 g/mol

g H2O= (1g/mL)(100mL)=100g
Q= (5.70 g CaCl2) / (110.9840 g CaCl2/mol) x (82.8 kJ/mol) = 4.2525 kJ = 4252.5 J

\Delta T= (4252.5 J) / [(4.184 J/g·°C)(100 g + 5.70 g)] = 9.62 °C change

23.0 °C + 9.62 °C = 32.62 °C